At the beginning of a walk, Roberto and juanita are 5.5 miles apart. If they leave at the same time and walk the same direction, Roberto overtakes juanita in 11 hours. If they walk towards each other they meet up in 1 hour. What are their speeds.
Answer by ankor@dixie-net.com(22740)

At the beginning of a walk, Roberto and juanita are 5.5 miles apart.
If they leave at the same time and walk the same direction, Roberto overtakes juanita in 11 hours.
If they walk towards each other they meet up in 1 hour.
What are their speeds.
:
let r = Robert's walking speed
let j = Juanita's walking speed
:
Write a distance equation for each scenario
:
Walk the same direction
11r - 11j = 5.5
simplify, divide by 11
r - j = .5
:
walk towards each other (for 1 hr)
r + j = 5.5
:
use elimination
r + j = 5.5
r - j = 0.5
---------------addition eliminates j, find r
2r = 6
r = 3 mph, Robert's speed
and
r + j = 5.5
3 + j = 5.5
j = 5.5 - 3
j = 2.5 mph is Juanita's speed